OpenStudy (sparklyme):

Looking for an explanation, not just an answer! Equation will be in comments. Please help!

OpenStudy (sparklyme):

Simplify. Assume that no variable equals zero. \[\frac{ (3x ^{-2}y ^{3})(5xy ^{-8}) }{ (x ^{-3})^{4}y ^{-2} }\]

OpenStudy (leenathan):

im busy but add the exponents

OpenStudy (leenathan):

nvm to much explaining let her answer

jhonyy9 (jhonyy9):

use this propertie of exponents x^a *x^b = x^(a+b) and x^a / x^b = x^(a-b) hope helped

OpenStudy (amorfide):

\[a^{m}a^{n}=a^{m+n}\] also remember multiplication is commutative, axb=bxa axbxc=axcxb etc so you would m ultiply the terms on the top of the fraction, then add the exponents

OpenStudy (mrs.ambrose614):

U use propriety of equations

OpenStudy (amorfide):

\[\frac{ a^{m} }{ a^{n} }=a^{m-n}\] @jhonyy9 beat me to it

jhonyy9 (jhonyy9):

np

jhonyy9 (jhonyy9):

yw

OpenStudy (sparklyme):

@steve816

OpenStudy (sparklyme):

@pooja195 @TheSmartOne